A 40 foot beam is to be cut in two. If one portion is one third the length of the other portion, what are the two length of the portions?
Answer by nerdybill(7384)

A 40 foot beam is to be cut in two. If one portion is one third the length of the other portion, what are the two length of the portions?
.
Let x = length of larger portion
then
x/3 = length of smaller portion
.
x + x/3 = 40
multiplying both sides by 3:
3x + x = 120
4x = 120
x = 30 feet (larger portion)
.
smaller portion:
x/3 = 30/3 = 10 feet